OpenStudy (barrelracer17):

Select interior, exterior, or on the circle (x-5)^2 +(y+3)^2=25 for the following point. (5, -3)

OpenStudy (misty1212):

guess i was not much help last time so lets just do it ok?

OpenStudy (barrelracer17):

Okay thank you and I'm sorry

OpenStudy (misty1212):

\[(x-5)^2+(y+3)^2=25\] looks like \[(x-h)^2+(y-k)^2=r^2\] you have to identify \(h, k, r\)

OpenStudy (misty1212):

this is an overlooked but important skill, recoginizing the specific from the general in this case \(h=5,k=-3\) and since \(r^2=25\) you know \(r=5\)

OpenStudy (misty1212):

that means your circle has center \((5,-3)\) and radius \(5\)

OpenStudy (misty1212):

since \((5,-3)\) is the center of the circle, it is certainly in the interior of the circle

OpenStudy (barrelracer17):

Ohhhh okay I get it now thank you so much

OpenStudy (misty1212):

hmm thanks for the compliment, but i seriously doubt you "get it" you will know if you get it when you can do one on your own try one

OpenStudy (barrelracer17):

Doing that right now :)

OpenStudy (barrelracer17):

I did it